# ULTHO: Ultra-Lightweight yet Efficient Hyperparameter Optimization in Deep Reinforcement Learning

### v1

@@ Abstract
Hyperparameter optimization (HPO) is a billion-dollar problem in machine learning, which significantly impacts the training efficiency and model performance. However, achieving efficient and robust HPO in deep reinforcement learning (RL) is consistently challenging due to its high non-stationarity and computational cost. To tackle this problem, existing approaches attempt to adapt common HPO techniques (e.g., population-based training or Bayesian optimization) to the RL scenario. However, they remain sample-inefficient and computationally expensive, which cannot facilitate a wide range of applications. In this paper, we propose ULTHO, an ultra-lightweight yet powerful framework for fast HPO in deep RL within single runs. Specifically, we formulate the HPO process as a multi-armed bandit with clustered arms (MABC) and link it directly to long-term return optimization. ULTHO also provides a quantified and statistical perspective to filter the HPs efficiently. We test ULTHO on benchmarks including ALE, Procgen, MiniGrid, and PyBullet. Extensive experiments demonstrate that the ULTHO can achieve superior performance with simple architecture, contributing to the development of advanced and automated RL systems.
@@ 1 Introduction
Deep reinforcement learning (RL) has propelled significant advancements across various fields, such as game playing [45, 54, 58], chip design [25], algorithm innovation [22, 42], and large language model (LLM) [46, 6, 12, 40]. However, training deep RL agents involves a number of design decisions and hyperparameter configurations, in which minor variations can result in substantial effects on learning efficiency and final performance [19]. While many efforts have been devoted to improving the design choices [27, 20, 28, 29] and implementation details [50, 3, 30, 65], RL-specific hyperparameter optimization (HPO) does not receive sufficient attention in the community. This gap is particularly notable given the high sensitivity of deep RL algorithms to HPs, which restricts their border applications and the development of advanced and automated RL systems [24].
@@
To tackle the HPO challenge in deep RL, several approaches have been proposed to dynamically adjust HPs during training. For instance, [32] introduced population-based training (PBT) that evolves a population of agents to explore and exploit different HP schedules, enabling adaptation to non-stationary settings. Notably, PBT is utilized to solve the complex 3D multiplayer game, Quake III Arena, achieving human-level performance by optimizing the agent policies along with internal reward signals in a hierarchical manner [33]. [24] further extended PBT and proposed SEARL, which optimizes the HPs and also the neural architecture while simultaneously training the agent, significantly improving sample efficiency by sharing experience across the population. However, PBT-like methods remain sample-inefficient and computationally expensive, limiting their practicality to a wider range of tasks. To address these issues, single-run HPO methods such as HOOF [49] have been developed. HOOF adapts HPs using off-policy importance sampling, optimizing a one-step improvement objective with sampled trajectories. As a gradient-free algorithm, HOOF can significantly reduce reduce overhead and promote sample efficiency. However, HOOF greedily focuses on HP configurations that maximize the value of the updated policy without introducing an exploration mechanism. Moreover, its reliance on importance sampling also makes it prone to high variance when policy distributions shift, hurting its robustness in highly dynamic environments.
@@ 2.1 HPO in Machine Learning
HPO is essential in ML as it significantly impacts model performance, convergence speed, and generalization ability. Prominent techniques, including random and grid search [11], Bayesian optimization (BO) [55, 60, 57], multi-fidelity search strategies [38, 21, 34], grdient-based methods [41, 61, 13, 44], population-based training [32, 47, 59], and RL-based methods [18, 35, 31]. For instance, [21] proposes BOHB that combines BO and Hyperband to balance the exploration-exploitation trade-off while dynamically allocating resources, significantly improving efficiency and robustness over a single method. [35] further extends PBT and proposed population-based bandits (PB2) by incorporating a MAB strategy to adaptively explore HP schedules, improving sample efficiency over standard PBT. [35] formulates HPO as a sequential decision problem and solves it with RL, which enables adaptive HP tuning and reduces reliance on hand-crafted acquisition functions. However, the methods above are prone to be computationally expensive and less effective for RL scenarios.
@@ 3.1 Reinforcement Learning
We study the RL problem considering a Markov decision process (MDP) [10, 36] defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,d_{0},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the extrinsic reward function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition function that defines a probability distribution over $\mathcal{S}$, $d_{0}\in\Delta(\mathcal{S})$ is the distribution of the initial observation $\bm{s}_{0}$, and $\gamma\in[0,1)$ is a discount factor. The goal of RL is to learn a policy $\pi_{\bm{\theta}}(\bm{a}|\bm{s})$ to maximize the expected discounted return:
@@ 4.1 MABC for HPO
2:  Initialize a set $\bm{\Psi}$ of HPs, an exploration coefficient $c$, a window length $W$ for estimating the Q-functions;
3:  $\forall\bm{\psi}\in\bm{\Psi},\forall\psi\in\bm{\psi}$, let
$N(\bm{\psi})=1,Q(\bm{\psi})=0,R(\bm{\psi})={\rm FIFO}(W)$
$N(\psi)=1,Q(\psi)=0,R(\psi)={\rm FIFO}(W)$
@@
6:     Perform the generalized advantage estimation (GAE) to get the estimated returns;
@@
10:     Compute the mean return $\bar{V}_{\bm{\phi}}$ obtained by the new policy;
11:     Add $\bar{V}_{\bm{\phi}}$ to the queue $R(\bm{\psi}_{e})$ and $R(\psi_{e})$ using the first-in-first-out rule;
12:     Update $Q(\bm{\psi}_{e})$ and $Q(\psi_{e})$ using Eq. (6);
@@
We solve the defined MABC problem by hierarchically applying the upper confidence bound (UCB) [4] algorithm, which is an effective and widely-used method to balance exploration-exploitation in bandit problems. Specifically, at each time step $t$, we first select a cluster from $\bm{\Psi}$ by the following policy:
$\bm{\psi}_{t}=\underset{\bm{\psi}\in\bm{\Psi}}{\rm argmax}\left[Q_{t}(\bm{\psi})+c\sqrt{\frac{\log t}{N_{t}(\bm{\psi})}}\right],$ (5)
where $N_{t}(\bm{\psi})$ is the number of times that $\bm{\psi}$ has been chosen before time step $t$, and $c$ is the exploration coefficient. Before the $t$-th update, we select a $\bm{\psi}$ using Eq. (5), which will used for the policy updates. Then, the counter is updated by $N_{t}(\bm{\psi})=N_{t}(\bm{\psi})+1$. Next, we collect rollouts with the new policy and update the Q-function using a sliding window average of the past mean returns obtained by the agent after being updated using $\bm{\psi}$:
$Q_{t}(\bm{\psi})=\frac{1}{W}\sum_{i=1}^{W}\bar{V}_{\bm{\phi}}[i],\bar{V}_{\bm{\phi}}=\frac{1}{T}\sum_{t=1}^{T}V_{\bm{\phi}}(\bm{s}_{t}),$ (6)
where $V_{\bm{\phi}}(\bm{s}_{t})$ is the estimated return predicted by the value network and $T$ is the episode length.
@@
$\psi_{t}=\underset{\psi\in\bm{\psi}_{t}}{\rm argmax}\left[Q_{t}(\psi)+c\sqrt{\frac{\log t}{N_{t}(\psi)}}\right],$ (7)
where the corresponding $Q$-function and counter are updated following the same way as Eq. (5).
@@
1:  Initialize a set $\bm{\Psi}$ of HP clusters and execute Algorithm 1;
@@ 4.2 Relay-ULTHO
Once the COI and NOI are identified, the optimization process is focused on these clusters, allowing the algorithm to refine the most influential cluster and, further explore the neglected cluster. This approach can improve the completeness of exploration, thereby squeezing out additional performance gains. However, we also need to highlight that this algorithm may result in sub-optimal performance due to its restricted landscape.
@@ 5 Experiments
  • Q1: Can ULTHO improve performance as compared to using fixed HP values? (See Figure 4, 5, 9, 10, 20, and Table 1)
  • Q2: Can the relay optimization further enhance the ULTHO algorithm? (See Figure 4, 5, 9, 10, 20, and Table 1)
  • Q3: What are the detailed decision processes of the ULTHO algorithm? (See Figure 6, 7, 21, and 22)
@@ 5.1.1 Benchmark Selection
We first evaluate the ULTHO using the arcade learning environment (ALE) benchmark [9], a collection of arcade game environments that requires the agent to learn motor control directly from images. Specifically, we focus on a subset of ALE known as ALE-5, which typically produces median score estimates for all games that are within 10% of their true values [2]. For all environments, we stack 4 consecutive frames to form an input state with the data shape of $(84,84,4)$. Additionally, we introduce the Procgen benchmark with 16 procedurally-generated environments. Procegn is similar to the ALE benchmark yet involves much higher dynamicity and presents a more difficult challenge for HPO. All the environments use a discrete fifteen-dimensional action space and generate $(64,64,3)$ RGB observations, and we use the easy mode and train the agents on 200 levels before testing them on the full distribution of levels. Furthermore, we introduce the MiniGrid [15] and PyBullet [17] to test ULTHO in sparse-rewards environments and continuous control tasks.
@@ 5.1.2 Algorithmic Baselines
Figure 4: Performance comparison of the two ULTHO algorithms and HPO baselines on the ALE-5 benchmark, in which the mean and standard error are computed using five random seeds. All the scores are normalized using the min-max normalization with the human expert performance. The two ULTHO algorithms achieve similar or even higher performance with a few training budgets as compared to other HPO baselines.
Figure 5: Training performance of the PPO and its combinations with two ULTHO algorithms in eight Procgen environments, in which the mean and standard deviation are computed using five random seeds. ULTHO consistently outperforms the vanilla PPO agent in all the environments, significantly improving the sample efficiency. However, the Relay-ULTHO does not achieve any further performance gains, indicating that a more comprehensive policy, which continuously adapts multiple HPs is crucial for highly dynamic environments.
@@ 5.1.4 Evaluation Metrics
For the baselines that require multiple training runs (e.g., PBT), we use the maximum observed past scores as their final performance. For our ULTHO algorithm, we simply use the observed episode return at the end of training as its performance. Note that the score of each method on each environment is computed as the average episode returns over 100 episodes and 5 random seeds. Additionally, we also compare the computational efficiency by evaluating the required training budgets of all the methods.
@@ 5.2.1 Comparison with HPO Baselines
We first compare the performance of ULTHO and four HPO baselines on the ALE benchmark, and their normalized scores and training budgets are illustrated in Figure 4. Here, the baseline PPO is trained using the common HPs reported in its origin paper [53]. ULTHO successfully outperforms the PPO baseline in all five environments and achieves the highest performance in two environments, especially in Q*Bert environment. In contrast, the SMAC and SMAC+HB rank second and third regarding the average performance across all the environments. The PBT method excels in the BattleZone environment, but it fails to outperform the baseline PPO in the Q*bert environment. By adaptively tuning HPs within the training run, ULTHO can achieve remarkable efficiency with much lower computational cost as compared to the baselines.
@@
Next, we report the performance of ULTHO on the Procgen benchmark. Similarly, the baseline PPO is trained using the reported HPs in [16]. Figure 5 illustrates the training performance comparison in eight environments, and the full curves are provided in Figure 20. ULTHO outperforms the baseline PPO in 15 environments, achieving significant performance gains in environments like BigFish, Chaser, and StarPilot. Additionally, we evaluate ULTHO on the full distribution of levels, and Table 1 illustrates the performance comparison with three HPO baselines. ULTHO outperforms these baselines regarding both the aggregated mean and IQM. These results demonstrate that ULTHO can effectively select the appropriate HPs at different learning stages and significantly improve learning efficiency even in highly dynamic environments.
@@ 5.2.3 The Detailed Decision Process
Furthermore, we analyze the detailed decision processes of ULTHO. Figure 6(a) illustrates the cumulative proportion of each cluster selected during the whole training. It is clear that ULTHO primarily selects the NUE cluster, while the other three clusters account for approximately 20%. For intra-cluster decisions, we find that the LR=5e-4, VLC=1.0, BS=2048, and NUE=2 are the most popular choices. Additionally, Figure 7 illustrates the variation of aggregated confidence intervals of these clusters during the training. ULTHO shows the explicitly distinct preferences of HP clusters, which aligns with the cumulative proportion illustrated in Figure 6(a). These confidence intervals provide a statistical and quantified perspective to identify the importance of each HP cluster, contributing to a more efficient filtering approach. Finally, we provide the detailed decision processes of each method and environment in Appendix D.

### v2

@@ Abstract
Hyperparameter optimization (HPO) is a billion-dollar problem in machine learning, which significantly impacts the training efficiency and model performance. However, achieving efficient and robust HPO in deep reinforcement learning (RL) is consistently challenging due to its high non-stationarity and computational cost. To tackle this problem, existing approaches attempt to adapt common HPO techniques (e.g., population-based training or Bayesian optimization) to the RL scenario. However, they remain sample-inefficient and computationally expensive, which cannot facilitate a wide range of applications. In this paper, we propose ULTHO, an ultra-lightweight yet powerful framework for fast HPO in deep RL within single runs. Specifically, we formulate the HPO process as a multi-armed bandit with clustered arms (MABC) and link it directly to long-term return optimization. ULTHO also provides a quantified and statistical perspective to filter the HPs efficiently. We test ULTHO on benchmarks including ALE, Procgen, MiniGrid, and PyBullet. Extensive experiments demonstrate that the ULTHO can achieve superior performance with a simple architecture, contributing to the development of advanced and automated RL systems. Our code is available at the GitHub repository [^1] https://github.com/yuanmingqi/ULTHO.
@@ 1 Introduction
Deep reinforcement learning (RL) has propelled significant advancements across various fields, such as game playing [46, 55, 60], chip design [25], algorithm innovation [22, 43], and large language model (LLM) [47, 6, 12, 41]. However, training deep RL agents involves a number of design decisions and hyperparameter configurations, in which minor variations can result in substantial effects on learning efficiency and final performance [19]. While many efforts have been devoted to improving the design choices [28, 20, 29, 30] and implementation details [51, 3, 31, 67], RL-specific hyperparameter optimization (HPO) does not receive sufficient attention in the community. This gap is particularly notable given the high sensitivity of deep RL algorithms to HPs, which restricts their broader applications and the development of advanced and automated RL systems [24].
@@
To tackle the HPO challenge in deep RL, several approaches have been proposed to dynamically adjust HPs during training. For instance, [33] introduced population-based training (PBT) that evolves a population of agents to explore and exploit different HP schedules, enabling adaptation to non-stationary settings. Notably, PBT is utilized to solve the complex 3D multiplayer game, Quake III Arena, achieving human-level performance by optimizing the agent policies along with internal reward signals in a hierarchical manner [34]. [24] further extended PBT and proposed SEARL, which optimizes the HPs and also the neural network architecture while simultaneously training the agent, significantly improving sample efficiency by sharing experience across the population. However, PBT-like methods remain sample-inefficient and computationally expensive, limiting their practicality to a wider range of tasks. To address these issues, single-run HPO methods such as HOOF [50] have been developed. HOOF adapts HPs using off-policy importance sampling, optimizing a one-step improvement objective with sampled trajectories. As a gradient-free algorithm, HOOF can significantly reduce overhead and promote sample efficiency. However, HOOF greedily focuses on HP configurations that maximize the value of the updated policy without introducing an exploration mechanism. Moreover, its reliance on importance sampling also makes it prone to high variance when policy distributions shift, hurting its robustness in highly dynamic environments.
@@ 2.1 HPO in Machine Learning
HPO is essential in machine learning as it significantly impacts model performance, convergence speed, and generalization ability. Prominent techniques, including random and grid search [11], Bayesian optimization (BO) [56, 62, 59], multi-fidelity search strategies [39, 21, 35], grdient-based methods [42, 63, 13, 45], population-based training [33, 48, 61], and RL-based methods [18, 36, 32]. For instance, [21] proposes BOHB that combines BO and Hyperband to balance the exploration-exploitation trade-off while dynamically allocating resources, significantly improving efficiency and robustness over a single method. [36] further extends PBT and proposed population-based bandits (PB2) by incorporating a MAB strategy to adaptively explore HP schedules, improving sample efficiency over standard PBT. [36] formulates HPO as a sequential decision problem and solves it with RL, which enables adaptive HP tuning and reduces reliance on hand-crafted acquisition functions. However, the methods above are prone to be computationally expensive and less effective for RL scenarios.
@@ 3.1 Reinforcement Learning
We study the RL problem considering a Markov decision process (MDP) [10, 37] defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,d_{0},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the extrinsic reward function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition function that defines a probability distribution over $\mathcal{S}$, $d_{0}\in\Delta(\mathcal{S})$ is the distribution of the initial observation $\bm{s}_{0}$, and $\gamma\in[0,1]$ is a discount factor. The goal of RL is to learn a policy $\pi_{\bm{\theta}}(\bm{a}|\bm{s})$ to maximize the expected discounted return:
@@ 4.1 MABC for HPO
2: Initialize a set $\bm{\Psi}$ of HPs, an exploration coefficient $c$, a window length $W$ for estimating the utility functions;
[…]
$N(\bm{\psi})=1,U(\bm{\psi})=0,R(\bm{\psi})={\rm FIFO}(W)$
$N(\psi)=1,U(\psi)=0,R(\psi)={\rm FIFO}(W)$
@@
6:  Perform the generalized advantage estimation (GAE) to get the estimated task returns;
@@
10:  Compute $\bar{V}_{\bm{\phi}}$ obtained by the new policy;
11:  Add $\bar{V}_{\bm{\phi}}$ to the queue $R(\bm{\psi}_{e})$ and $R(\psi_{e})$;
12:  Update $U(\bm{\psi}_{e})$ and $U(\psi_{e})$ using Eq. (7);
@@
We solve the defined MABC problem by hierarchically applying the upper confidence bound (UCB) [4] algorithm, which is an effective and widely-used method to balance exploration-exploitation in bandit problems. Specifically, at the $i$-th episode, we first select a cluster from $\bm{\Psi}$ by the following policy:
$\underbrace{\bm{\psi}_{i}=\underset{\bm{\psi}\in\bm{\Psi}}{\rm argmax}\left[U_{i}(\bm{\psi})+c\sqrt{\frac{\log i}{N_{i}(\bm{\psi})}}\right]}_{\text{Inter-cluster Scheduling Policy}},$ (5)
where $N_{i}(\bm{\psi})$ is the number of times that $\bm{\psi}$ has been chosen before the $i$-the episode, and $c$ is the exploration coefficient.
@@
$\underbrace{\psi_{i}=\underset{\psi\in\bm{\psi}_{i}}{\rm argmax}\left[U_{i}(\psi)+c\sqrt{\frac{\log i}{N_{i}(\psi)}}\right]}_{\text{Intra-cluster Scheduling Policy}}.$ (6)
The selected HP value will be used for policy updates. Then the utility functions of the cluster and the corresponding value are updated using a sliding window average of the past mean episode returns obtained by the agent:
$\underbrace{U_{i}(\bm{\psi})=\frac{1}{W}\sum_{j=1}^{W}\bar{V}_{\bm{\phi}}[j],\bar{V}_{\bm{\phi}}=\frac{1}{T}\sum_{t=1}^{T}V_{\bm{\phi}}(\bm{s}_{t})}_{\text{Utility Evaluation}},$ (7)
where $V_{\bm{\phi}}(\bm{s}_{t})$ is the estimated return predicted by the value network, $T$ is the episode length, and $W$ is the window length.
@@
1: Execute the Algorithm 1;
@@ 4.2 Relay-ULTHO
Once the COI and NOI are identified, the optimization process is focused on these clusters, allowing the algorithm to refine the most influential cluster and further explore the neglected cluster. This approach can improve the completeness of exploration, thereby squeezing out additional performance gains. However, we also need to highlight that this algorithm may result in sub-optimal performance due to its restricted landscape.
@@ 5 Experiments
  • Q1: Can ULTHO improve performance as compared to using fixed HP values? (See Figure 4, 5, 9, 10, 20, 21, and Table 1)
  • Q2: Can the relay optimization further enhance the ULTHO algorithm? (See Figure 4, 5, 9, 10, 20, 21, and Table 1)
  • Q3: What are the detailed decision processes of the ULTHO algorithm? (See Figure 6, 7, 24, and 25)
@@ 5.1.1 Benchmark Selection
We first evaluate the ULTHO using the arcade learning environment (ALE) benchmark [9], a collection of arcade game environments that requires the agent to learn motor control directly from images. Specifically, we focus on a subset of ALE known as ALE-5, which typically produces median score estimates for all games that are within 10% of their true values [2]. For all environments, we stack 4 consecutive frames to form an input state with the data shape of (84,84,4)(84,84,4)( 84 , 84 , 4 ). Additionally, we introduce the Procgen benchmark with 16 procedurally-generated environments. Procegn is similar to the ALE benchmark yet involves much higher dynamicity and presents a more difficult challenge for HPO. All the environments use a discrete fifteen-dimensional action space and generate (64,64,3)(64,64,3)( 64 , 64 , 3 ) RGB observations, and we use the easy mode and train the agents on 200 levels before testing them on the full distribution of levels. Furthermore, we introduce the MiniGrid [15] and PyBullet [17] to test ULTHO in sparse-rewards environments and continuous control tasks.
@@ 5.1.2 Algorithmic Baselines
Figure 4: Performance comparison of the two ULTHO algorithms and HPO baselines on the ALE-5 benchmark, in which the mean and standard error are computed using five random seeds. All the scores are normalized using the min-max normalization with the human expert performance. The two ULTHO algorithms achieve similar or even higher performance with a few training budgets as compared to other HPO baselines. An ablation study that uses the same training budgets for all the methods can be found in Appendix D.
Figure 5: Training performance of the PPO and its combinations with two ULTHO algorithms in eight Procgen environments, in which the mean and standard deviation are computed using five random seeds. ULTHO consistently outperforms the vanilla PPO agent in all the environments, significantly improving the sample efficiency. However, the Relay-ULTHO does not achieve any further performance gains, indicating that a more comprehensive policy, which continuously adapts multiple HPs, is crucial for highly dynamic environments.
@@ 5.1.4 Evaluation Metrics
For the baselines that require multiple training runs (e.g., PBT), we use the maximum observed past scores as their final performance. For our ULTHO algorithm, we simply use the observed episode return at the end of training as its performance. Note that the score of each method on each environment is computed as the average episode returns over 100 episodes and 5 random seeds. Additionally, we compare computational efficiency by evaluating the required training budgets for all the methods.
@@ 5.2.1 Comparison with HPO Baselines
We first compare the performance of ULTHO and four HPO baselines on the ALE benchmark, and their normalized scores and training budgets are illustrated in Figure 4. Here, the baseline PPO is trained using the common HPs reported in its original paper [54]. ULTHO successfully outperforms the PPO baseline in all five environments and achieves the highest performance in two environments, especially in Q*Bert environment. In contrast, the SMAC and SMAC+HB rank second and third in terms of average performance across all the environments. The PBT method excels in the BattleZone environment, but it fails to outperform the baseline PPO in the Q*bert environment. By adaptively tuning HPs within the training run, ULTHO can achieve remarkable efficiency with much lower computational cost as compared to the baselines.
@@
Next, we report the performance of the ULTHO on the Procgen benchmark. Similarly, the baseline PPO is trained using the reported HPs in [16]. Figure 5 illustrates the training performance comparison in eight environments, and the full curves are provided in Figure 20. ULTHO outperforms the baseline PPO in 15 environments, achieving significant performance gains in environments like BigFish, Chaser, and StarPilot. Additionally, we evaluate ULTHO on the full distribution of levels, and Table 1 illustrates the performance comparison with three HPO baselines. ULTHO outperforms these baselines regarding both the aggregated mean and IQM. These results demonstrate that ULTHO can effectively select the appropriate HPs at different learning stages and significantly improve learning efficiency even in highly dynamic environments.
@@ 5.2.3 The Detailed Decision Process
Furthermore, we analyze the detailed decision processes of ULTHO. Figure 6(a) illustrates the cumulative proportion of each cluster selected during the whole training. It is clear that ULTHO primarily selects the NUE cluster, while the other three clusters account for approximately 20%. For intra-cluster decisions, we find that the LR=5e-4, VLC=1.0, BS=2048, and NUE=2 are the most popular choices. Additionally, Figure 7 illustrates the variation of aggregated confidence intervals of these clusters during the training. ULTHO shows the explicitly distinct preferences of HP clusters, which aligns with the cumulative proportion illustrated in Figure 6(a). These confidence intervals provide a statistical and quantified perspective to identify the importance of each HP cluster, contributing to a more efficient filtering approach. Finally, we provide the detailed decision processes of each method and environment in Appendix E.